OpenStudy (calculusxy):

Can someone explain to me how to work with the slope-intercept formula: y=mx+b ?

OpenStudy (campbell_st):

so what is it about the slope intercept equation that causes you a problem

OpenStudy (calculusxy):

I need help on understanding the different variables that the formula has. @campbell_st

OpenStudy (campbell_st):

ok... so x is the independent and y is the dependent variables.... m is the slope of the line b is where the line cuts the y-axis... called the y-intercept so y = 3x + 5 the slope is 3 and the y-intercept is 5 or (0, 5) y = 10 - 2x the slope m = -2 and y-intercept is 10 y = 12 slope = 0 y-intercept is 12 hope it helps

OpenStudy (kc_kennylau):

That means for any (x,y) pair, y=mx+b the equality holds.